use hn the unemployment rate O An increase in the population OA technological breakthrough. QUESTION 3 freeze in Peru causes the price of coffeeto skhyrocket. Assuming coffee and tea are substitutes, which of the following wil happen? The quantity demanded of coffee will decrease, and the demand for tea will increase The demand for coffee will increase, and the demand for tea will increase The demand for coffee will decrease, and the quantity demanded of tea will increase O The quantity demanded of coffee will increase, and the quantity demanded of tea will increase. The quantity demanded of coffee will increase, and the demand for tea will increase. QUESTION4 Combination CltingFoo 110 105 10 15 20 25 30 35 additional unit of food is approximately According to the table above, if the economy is currently producing at point C the opportunity cost of one O 107 units of clothing s units of clothing 1/3 units of clothing 1/2 units of clothing 1/4 units of clothing Click Sove and Submit to sove and submit. Cick Sove All Answers to sove ad answrs
Answer #1

3. Ans: The quantity demanded of coffee will decrease, and the demand for tea will increase.

Explanation:

When price of coffee increased, according to law of demand, the quantity demanded of coffee will decrease. And some of consumer now switch to consume tea instead of coffee. So, the demand for tea will increase.

4. Ans: 1 / 2 units of clothing

Explanation:

The opportunity cost of one additional unit of food = (10 - 5) / (105 - 95) = 5 / 10 = 1 / 2 units of clothing.